# A system of NLS arising in optical material without Galilean symmetry

### v1

@@ 1 Introduction and Main Results
1. Using mass, energy conservation and the Gagliardo-Nirenberg inequality, one can easily obtain the global existence with the initial data $(u_{0},v_{0})\in H^{1}(\mathbb{R})\times H^{1}(\mathbb{R})$ in one dimensional case. For three dimensional case, we only obtain the global existence in the set $A_{\omega,c}^{+}$, see (4.1) and Lemma 4.2, this is similar to the classical case.
2. From [18], the condition (1.15) is sharp. But in our result, we give a new condition: (1.19) and the restriction one of the initial data $\|u_{0}\|_{L^{2}}$ or $\|v_{0}\|_{L^{2}}$, for the existence of global solution. This condition is weaker than (1.15) since $M_{3\gamma}(Q,P)<M_{3\gamma}(Q^{*},P^{*})$. This new phenomenon means that when $\gamma\neq 3$, the momentum change of the initial data by (1.20) essentially influences the global properties of the solution, which comes from the lack of Galilean invariance.
3. This is different from the global existence in [9, Theorem 1.3], where they only restrict one of the $u(0)$ or $v(0)$. In fact, in [9], they considered the nonlinear NLS system with quadratic interaction, but in our case, we must add the condition (1.19), which comes from the cubic interaction.
This paper is organized as follows:. In Section 2, we study the variational problem and show the existence and non-existence of traveling solitary wave solutions (Theorem 1.1 and 1.2). In Section 3, we establish the high frequency limit of the traveling solitary waves. In Section 4, we establish the global existence (Theorem 1.5). The final section is an appendix.
2 Existence and non-existence of traveling solitary waves
@@
$Q_{\omega,c}(u,v)=$ $\frac{1}{2}\int|\nabla u|^{2}+\frac{1}{2}\int|\nabla v|^{2}+\frac{\omega}{2}\|u\|_{L^{2}}^{2}+\frac{3\omega\gamma}{2}\|v\|_{L^{2}}^{2}+\frac{1}{2}c\cdot P(u,v).$ (2.1)
The energy functional (1.10) reads
$S_{\omega,c}(u,v)=Q_{\omega,c}(u,v)-D(u,v),$
@@
$N_{\omega,c}(u,v)=\partial_{\tau}S_{\omega,c}(\tau u,\tau v)|_{\tau=1}=2Q_{\omega,c}(u,v)-4D(u,v).$ (2.2)
@@
$\mu_{\omega,c}=\inf\{S_{\omega,c}(\phi,\psi)~{}:~{}(\phi,\psi)\in\mathcal{N}_{\omega,c}\},$ (2.3)
@@
$\mathcal{N}_{\omega,c}=\{(\phi,\psi)\in X_{\omega,c},~{}(\phi,\psi)\neq(0,0),~{}N_{\omega,c}(\phi,\psi)=0\}.$
We define the minimizers $a_{\omega,c}$ by
$a_{\omega,c}=\{(\phi,\psi)\in\mathcal{N}_{\omega,c}~{}:~{}S_{\omega,c}(\phi,\psi)=\mu_{\omega,c}\}.$
@@
$\widetilde{Q}_{\omega,c}(u,v)=$ $\frac{1}{2}\int|\nabla u|^{2}+\frac{1}{2}\int|\nabla v|^{2}+\frac{1}{2}\left(\omega-\frac{|c|^{2}}{4}\right)\|u\|_{L^{2}}^{2}+\frac{1}{2}\left(3\gamma\omega-\frac{\gamma^{2}|c|^{2}}{4}\right)\|v\|_{L^{2}}^{2},$
@@
$\tilde{Q}_{\omega,c}(\tilde{u},\tilde{v})=Q_{\omega,c}(u,v),~{}~{}\widetilde{D}(\tilde{u},\tilde{v})=D(u,v).$
@@
$\tilde{S}_{\omega,c}(u,v)=$ $\tilde{Q}_{\omega,c}(u,v)-\tilde{D}(u,v),$
$\tilde{N}_{\omega,c}(u,v)=$ $\partial_{\tau}\tilde{S}_{\omega,c}(\tau u,\tau v)|_{\tau=1}=2\tilde{Q}_{\omega,c}(u,v)-4\tilde{D}(u,v).$
@@
$\widetilde{\mathcal{A}}_{\omega,c}=\{(\phi,\psi)\in\tilde{X}_{\omega,c}~{}:~{}(\phi,\psi)\neq(0,0),~{}\tilde{S}^{\prime}_{\omega,c}(\phi,\psi)=0\},$
$\widetilde{\mathcal{G}}_{\omega,c}=\{(\phi,\psi)\in\widetilde{\mathcal{A}}_{\omega,c},~{}\tilde{S}_{\omega,c}(\phi,\psi)\leq\tilde{S}_{\omega,c}(\phi_{1},\psi_{1}),~{}\text{for all}~{}(\phi_{1},\psi_{1})\in\tilde{\mathcal{A}}_{\omega,c}\},$
$\tilde{\mu}_{\omega,c}=\inf\{\tilde{S}_{\omega,c}(\phi,\psi)~{}:~{}(\phi,\psi)\in\tilde{\mathcal{N}}_{\omega,c}\},$
$\tilde{\mathcal{N}}_{\omega,c}=\{(\phi,\psi)\in\tilde{X}_{\omega,c},~{}(\phi,\psi)\neq(0,0),~{}\tilde{N}_{\omega,c}(\phi,\psi)=0\},$
$\tilde{a}_{\omega,c}=\{(\phi,\psi)\in\tilde{\mathcal{N}}_{\omega,c}~{}:~{}\tilde{S}_{\omega,c}(\phi,\psi)=\tilde{\mu}_{\omega,c}\}.$
@@
$(\phi,\psi)\in{\mathcal{G}}_{\omega,c}\Leftrightarrow(\tilde{\phi},\tilde{\psi})\in\tilde{\mathcal{G}}_{\omega,c},~{}~{}(\phi,\psi)\in a_{\omega,c}\Leftrightarrow(\tilde{\phi},\tilde{\psi})\in\tilde{a}_{\omega,c},$
@@
$\mu_{\omega,c}=\tilde{\mu}_{\omega,c}.$
@@
$\int\left(|\nabla\phi|^{2}+|\nabla\psi|^{2}\right)+\omega\int\left(|\phi|^{2}+3\gamma|\psi|^{2}\right)+(ic\cdot\nabla\phi,\phi)+(i\gamma c\cdot\nabla\psi,\psi)=4D(\phi,\psi).$ (2.4)
@@
$\frac{N-2}{2}\int\left(|\nabla\phi|^{2}+|\nabla\psi|^{2}\right)+\frac{N}{2}\int\left(\omega|\phi|^{2}+3\gamma\omega|\psi|^{2}\right)$
@@
$\mathcal{G}_{\omega,c}=a_{\omega,c}\neq\emptyset.$
In fact, this is equivalent to $\widetilde{\mathcal{G}}_{\omega,c}=\tilde{a}_{\omega,c}\neq\emptyset.$
Remark 2.3.
If $(u,v)\in\mathcal{G}_{\omega,c}$, then $v\neq 0$. In fact, if $v=0$, from system (1.9), one can easily obtain that $u=0$.
@@
Let $\gamma\neq 3$, then $\tilde{a}_{\omega,c}\subset\widetilde{\mathcal{G}}_{\omega,c}$.
@@
Let $(\phi,\psi)\in\tilde{a}_{\omega,c}$. Since $\tilde{N}_{\omega,c}(\phi,\psi)=0$ and $(\phi,\psi)\neq 0$, we have
$\left(\tilde{N}^{\prime}_{\omega,c}(\phi,\psi),(\phi,\psi)\right)=4\tilde{Q}_{\omega,c}(\phi,\psi)-16\tilde{D}(\phi,\psi)=-4\tilde{Q}_{\omega,c}(\phi,\psi)<0.$ (2.6)
By the Lagrange multiplier theorem, there exists $\lambda\in\mathbb{R}$ such that $\tilde{S}^{\prime}_{\omega,c}(\phi,\psi)=\lambda\tilde{N}^{\prime}_{\omega,c}(\phi,\psi)$. Moreover, we have
$\lambda\left(\tilde{N}^{\prime}_{\omega,c}(\phi,\psi),(\phi,\psi)\right)=\left(\tilde{S}^{\prime}_{\omega,c}(\phi,\psi),(\phi,\psi)\right)=\tilde{N}_{\omega,c}(\phi,\psi)=0.$ (2.7)
Combining (2.6) and (2.7), we can obtain $\lambda=0$. Hence, $\tilde{S}^{\prime}_{\omega,c}(\phi,\psi)=0$, which implies that $(\phi,\psi)\in\tilde{\mathcal{A}}_{\omega,c}$.
Notice that $\tilde{\mathcal{A}}_{\omega,c}\subset\tilde{\mathcal{N}}_{\omega,c}$ and $(\phi,\psi)\in\tilde{a}_{\omega,c}$, we have
$\tilde{S}_{\omega,c}(\phi,\psi)\leq\tilde{S}_{\omega,c}(\phi_{1},\psi_{1})~{}~{}\text{for all}~{}~{}(\phi_{1},\psi_{1})\in\tilde{\mathcal{A}}_{\omega,c},$
which implies $(\phi,\psi)\in\tilde{\mathcal{G}}_{\omega,c}$. ∎
@@
Let $1\leq N\leq 3$, $\gamma>0$ and $\gamma\neq 3$. Assume that $\tilde{a}_{\omega,c}\neq\emptyset$, then $\widetilde{\mathcal{G}}_{\omega,c}\subset\tilde{a}_{\omega,c}$.
@@
Let $(\phi,\psi)\in\widetilde{\mathcal{G}}_{\omega,c}$. One can take $(\phi_{1},\psi_{1})\in\tilde{a}_{\omega,c}$, by Lemma 2.4, we have $(\phi_{1},\psi_{1})\in\widetilde{\mathcal{G}}_{\omega,c}$, i.e., $\tilde{S}_{\omega,c}(\phi_{1},\psi_{1})=\tilde{S}_{\omega,c}(\phi,\psi)$. Therefore, for each $(u,v)\in\tilde{\mathcal{N}}_{\omega,c}$, we obtain
$\tilde{S}_{\omega,c}(\phi,\psi)=\tilde{S}_{\omega,c}(\phi_{1},\psi_{1})\leq\tilde{S}_{\omega,c}(u,v).$
Since $(\phi,\psi)\in\widetilde{\mathcal{G}}_{\omega,c}\subset\tilde{\mathcal{N}}_{\omega,c}$, we deduce that $(\phi,\psi)\in\tilde{a}_{\omega,c}$. ∎
@@
Let $1\leq N\leq 3$, $\gamma>0$ and $\gamma\neq 3$. Then $\tilde{\mu}_{\omega,c}>0$.
@@
By the definition of $\tilde{S}_{\omega,c}$, $\tilde{Q}_{\omega,c}$ and $\tilde{N}_{\omega,c}$, we have the following relation
$\tilde{S}_{\omega,c}=\frac{1}{2}\tilde{Q}_{\omega,c}+\frac{1}{4}\tilde{N}_{\omega,c}.$
One can rewrite $\tilde{\mu}_{\omega,c}$ as
$\tilde{\mu}_{\omega,c}=\inf\left\{\frac{1}{2}\tilde{Q}_{\omega,c},~{}~{}(\phi,\psi)\in\tilde{\mathcal{N}}_{\omega,c}\right\}.$ (2.8)
By Hölder inequality, Gagliardo-Nirenberg inequality and Sobolev embedding theorem, we have
$\tilde{Q}_{\omega,c}(u,v)=2\widetilde{D}(u,v)$
@@
$\leq$ $C\|\nabla u\|_{L^{2}}^{N}\|u\|_{L^{2}}^{4-N}+C\|\nabla v\|_{L^{2}}^{N}\|v\|_{L^{2}}^{4-N}+C\|\nabla u\|_{L^{2}}^{\frac{3N}{4}}\|u\|_{L^{2}}^{\frac{3(4-N)}{4}}\|\nabla v\|_{L^{2}}^{\frac{N}{4}}\|v\|_{L^{2}}^{\frac{(4-N)}{4}}$
$\leq$ $C\tilde{Q}^{2}_{\omega,c}(u,v).$
Dividing both sides of the above by $\tilde{Q}_{\omega,c}>0$, we obtain the positive lower bound for $\tilde{Q}_{\omega,c}$. ∎
@@
Let $1\leq N\leq 3$, $\gamma>0$ and $\gamma\neq 3$. If $(u,v)\in\tilde{X}_{\omega,c}$ satisfies $\tilde{N}_{\omega,c}(u,v)<0$, then $\frac{1}{2}\tilde{Q}_{\omega,c}>\tilde{\mu}_{\omega,c}$.
@@
If $\tilde{N}_{\omega,c}(u,v)<0$, then $02\tilde{D}(u,v)>\tilde{Q}_{\omega,c}(u,v)>$ 0. From this, we have
$\lambda_{0}:=\frac{\tilde{Q}_{\omega,c}(u,v)}{2\tilde{D}(u,v)}\in(0,1)$
and $\tilde{N}_{\omega,c}(\lambda_{0}u,\lambda_{0}v)=0$. By (2.8), we deduce
$\tilde{\mu}_{\omega,c}\leq\frac{1}{2}\tilde{Q}_{\omega,c}(\lambda_{0}u,\lambda_{0}v)=\frac{\lambda_{0}^{2}}{2}\tilde{Q}_{\omega,c}(u,v)<\frac{1}{2}\tilde{Q}_{\omega,c}(u,v).$
@@
Let $1\leq N\leq 3$, $\gamma>0$ and $\gamma\neq 3$. If the sequence $\{(u_{n},v_{n})\}$ weakly converges to $(u,v)\in\tilde{X}_{\omega,c}$, then
[…]
@@
Note that $\tilde{Q}_{\omega,c}$ and $\tilde{D}$ are invariant under
@@
$\tilde{Q}_{\omega,c}(\tilde{\tau}_{y}(u,v))=\tilde{Q}_{\omega,c}(u,v),~{}~{}\tilde{D}(\tilde{\tau}_{y}(u,v))=\tilde{D}(u,v),$

### v2

@@ 1 Introduction and Main Results
1. Using mass, energy conservation and the Gagliardo-Nirenberg inequality, one can easily obtain the global existence with the initial data $(u_{0,c},v_{0,c})\in H^{1}(\mathbb{R})\times H^{1}(\mathbb{R})$ in one-dimensional case. For three-dimensional case, we only obtain the global existence in the set $A_{\omega,c,\alpha-3\gamma}^{+}$, see (4.1) and Lemma 4.2, this is similar to the classical case.
2. From [17], the condition (1.15) is sharp with $\alpha=3\gamma=9$. However, when $\gamma\neq 3$ and $\alpha=3\gamma$, we give a new condition: (1.19) and the restriction one of the initial data $\|u_{0,c}\|_{L^{2}}$ or $\|v_{0,c}\|_{L^{2}}$, for the existence of a global solution. This condition is weaker than (1.15) since $M_{3\gamma}(Q,P)<M_{3\gamma}(Q^{*},P^{*})$. This new phenomenon means that when $\gamma\neq 3$, the momentum change of the initial data by (1.20) essentially influences the global properties of the solution, which comes from the lack of Galilean invariance.
3. This is different from the global existence in [8, Theorem 1.3], where they only restrict one of the $L^{2}$-norm of the initial data $u(0)$ or $v(0)$. In fact, in [8], they considered the nonlinear NLS system with quadratic interaction, but in our case, we must add the condition (1.19), which comes from the cubic interaction.
This paper is organized as follows: In Section 2, we study the variational problem and show the existence and non-existence of traveling solitary wave solutions (Theorem 1.1 and 1.2). In Section 3, we establish the high frequency limit of the traveling solitary waves. In Section 4, we establish the global existence (Theorem 1.5). The final section is an appendix.
2 Existence and non-existence of traveling waves
@@
$\mathcal{Q}_{\omega,c,\alpha-3\gamma}(u,v)=$ $\frac{1}{2}\int|\nabla u|^{2}+\frac{1}{2}\int|\nabla v|^{2}+\frac{\omega}{2}(\|u\|_{L^{2}}^{2}+3\gamma\|v\|_{L^{2}}^{2})+\frac{1}{2}(\alpha-3\gamma)\|v\|_{L^{2}}^{2}+\frac{1}{2}c\cdot P(u,v).$ (2.1)
The energy functional (1) reads
$S_{\omega,c,\alpha-3\gamma}(u,v)=\mathcal{Q}_{\omega,c,\alpha-3\gamma}(u,v)-D(u,v),$
@@
$N_{\omega,c,\alpha-3\gamma}(u,v)=\partial_{\tau}S_{\omega,c,\alpha-3\gamma}(\tau u,\tau v)|_{\tau=1}=2\mathcal{Q}_{\omega,c,\alpha-3\gamma}(u,v)-4D(u,v).$ (2.2)
@@
$\mu_{\omega,c,\alpha-3\gamma}=\inf\{S_{\omega,c,\alpha-3\gamma}(\phi,\psi)~{}:~{}(\phi,\psi)\in\mathcal{N}_{\omega,c,\alpha-3\gamma}\},$ (2.3)
@@
$\mathcal{N}_{\omega,c,\alpha-3\gamma}=\{(\phi,\psi)\in X_{\omega,c,\alpha-3\gamma}:~{}(\phi,\psi)\neq(0,0),~{}N_{\omega,c,\alpha-3\gamma}(\phi,\psi)=0\}.$
We define the minimizers $a_{\omega,c,\alpha-3\gamma}$ by
$a_{\omega,c,\alpha-3\gamma}=\{(\phi,\psi)\in\mathcal{N}_{\omega,c,\alpha-3\gamma}~{}:~{}S_{\omega,c,\alpha-3\gamma}(\phi,\psi)=\mu_{\omega,c,\alpha-3\gamma}\}.$
@@
$\widetilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}(u,v)=$ $\frac{1}{2}\int|\nabla u|^{2}+\frac{1}{2}\int|\nabla v|^{2}+\frac{1}{2}\left(\omega-\frac{|c|^{2}}{4}\right)\|u\|_{L^{2}}^{2}$
$+\frac{1}{2}\left(3\gamma\omega-\frac{\gamma^{2}|c|^{2}}{4}\right)\|v\|_{L^{2}}^{2}+\frac{\alpha-3\gamma}{2}\|v\|_{L^{2}}^{2},$
@@
$\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}(\tilde{u},\tilde{v})=\mathcal{Q}_{\omega,c,\alpha-3\gamma}(u,v),~{}~{}\widetilde{D}(\tilde{u},\tilde{v})=D(u,v).$
@@
$\tilde{S}_{\omega,c,\alpha-3\gamma}(u,v)=$ $\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}(u,v)-\tilde{D}(u,v),$
$\tilde{N}_{\omega,c,\alpha-3\gamma}(u,v)=$ $\partial_{\tau}\tilde{S}_{\omega,c,\alpha-3\gamma}(\tau u,\tau v)|_{\tau=1}=2\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}(u,v)-4\tilde{D}(u,v).$
@@
$\widetilde{\mathcal{A}}_{\omega,c,\alpha-3\gamma}=\{(\phi,\psi)\in\tilde{X}_{\omega,c,\alpha-3\gamma}:~{}(\phi,\psi)\neq(0,0),~{}\tilde{S}^{\prime}_{\omega,c,\alpha-3\gamma}(\phi,\psi)=0\},$
$\widetilde{\mathcal{G}}_{\omega,c,\alpha-3\gamma}=\{(\phi,\psi)\in\widetilde{\mathcal{A}}_{\omega,c,\alpha-3\gamma}:~{}\tilde{S}_{\omega,c,\alpha-3\gamma}(\phi,\psi)\leq\tilde{S}_{\omega,c,\alpha-3\gamma}(\phi_{1},\psi_{1}),~{}\text{for all}~{}(\phi_{1},\psi_{1})\in\tilde{\mathcal{A}}_{\omega,c,\alpha-3\gamma}\},$
$\tilde{\mu}_{\omega,c,\alpha-3\gamma}=\inf\{\tilde{S}_{\omega,c,\alpha-3\gamma}(\phi,\psi)~{}:~{}(\phi,\psi)\in\tilde{\mathcal{N}}_{\omega,c,\alpha-3\gamma}\},$
$\tilde{\mathcal{N}}_{\omega,c,\alpha-3\gamma}=\{(\phi,\psi)\in\tilde{X}_{\omega,c,\alpha-3\gamma}:~{}(\phi,\psi)\neq(0,0),~{}\tilde{N}_{\omega,c,\alpha-3\gamma}(\phi,\psi)=0\},$
$\tilde{a}_{\omega,c,\alpha-3\gamma}=\{(\phi,\psi)\in\tilde{\mathcal{N}}_{\omega,c,\alpha-3\gamma}:~{}\tilde{S}_{\omega,c,\alpha-3\gamma}(\phi,\psi)=\tilde{\mu}_{\omega,c,\alpha-3\gamma}\}.$
@@
$(\phi,\psi)\in{\mathcal{G}}_{\omega,c,\alpha-3\gamma}\Leftrightarrow(\tilde{\phi},\tilde{\psi})\in\tilde{\mathcal{G}}_{\omega,c,\alpha-3\gamma},~{}~{}(\phi,\psi)\in a_{\omega,c,\alpha-3\gamma}\Leftrightarrow(\tilde{\phi},\tilde{\psi})\in\tilde{a}_{\omega,c,\alpha-3\gamma},$
@@
$\mu_{\omega,c,\alpha-3\gamma}=\tilde{\mu}_{\omega,c,\alpha-3\gamma}.$
@@
$\int\left(|\nabla\phi|^{2}+|\nabla\psi|^{2}\right)$ $+\omega\int\left(|\phi|^{2}+3\gamma|\psi|^{2}\right)+(\alpha-3\gamma)\int|\psi|^{2}$
$+(ic\cdot\nabla\phi,\phi)+(i\gamma c\cdot\nabla\psi,\psi)=4D(\phi,\psi),$ (2.4)
@@
$\frac{N-2}{2}\int\left(|\nabla\phi|^{2}+|\nabla\psi|^{2}\right)+\frac{N}{2}\int\left(\omega|\phi|^{2}+3\gamma\omega|\psi|^{2}\right)+\frac{N}{2}(\alpha-3\gamma)\int|\psi|^{2}$
@@
$\mathcal{G}_{\omega,c,\alpha-3\gamma}=a_{\omega,c,\alpha-3\gamma}\neq\emptyset.$
In fact, this is equivalent to $\widetilde{\mathcal{G}}_{\omega,c,\alpha-3\gamma}=\tilde{a}_{\omega,c,\alpha-3\gamma}\neq\emptyset.$
@@
Let $\gamma\neq 3$, then $\tilde{a}_{\omega,c,\alpha-3\gamma}\subset\widetilde{\mathcal{G}}_{\omega,c,\alpha-3\gamma}$.
@@
Let $(\phi,\psi)\in\tilde{a}_{\omega,c,\alpha-3\gamma}$. Since $\tilde{N}_{\omega,c,\alpha-3\gamma}(\phi,\psi)=0$ and $(\phi,\psi)\neq 0$, we have
$\left(\tilde{N}^{\prime}_{\omega,c,\alpha-3\gamma}(\phi,\psi),(\phi,\psi)\right)=4\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}(\phi,\psi)-16\tilde{D}(\phi,\psi)=-4\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}(\phi,\psi)<0.$ (2.6)
By the Lagrange multiplier theorem, there exists $\lambda\in\mathbb{R}$ such that $\tilde{S}^{\prime}_{\omega,c,\alpha-3\gamma}(\phi,\psi)=\lambda\tilde{N}^{\prime}_{\omega,c,\alpha-3\gamma}(\phi,\psi)$. Moreover, we have
$\lambda\left(\tilde{N}^{\prime}_{\omega,c,\alpha-3\gamma}(\phi,\psi),(\phi,\psi)\right)=\left(\tilde{S}^{\prime}_{\omega,c,\alpha-3\gamma}(\phi,\psi),(\phi,\psi)\right)=\tilde{N}_{\omega,c,\alpha-3\gamma}(\phi,\psi)=0.$ (2.7)
Combining (2.6) and (2.7), we can obtain $\lambda=0$. Hence, $\tilde{S}^{\prime}_{\omega,c,\alpha-3\gamma}(\phi,\psi)=0$, which implies that $(\phi,\psi)\in\tilde{\mathcal{A}}_{\omega,c,\alpha-3\gamma}$.
Notice that $\tilde{\mathcal{A}}_{\omega,c,\alpha-3\gamma}\subset\tilde{\mathcal{N}}_{\omega,c,\alpha-3\gamma}$ and $(\phi,\psi)\in\tilde{a}_{\omega,c,\alpha-3\gamma}$, we have
$\tilde{S}_{\omega,c,\alpha-3\gamma}(\phi,\psi)\leq\tilde{S}_{\omega,c,\alpha-3\gamma}(\phi_{1},\psi_{1})~{}~{}\text{for all}~{}~{}(\phi_{1},\psi_{1})\in\tilde{\mathcal{A}}_{\omega,c,\alpha-3\gamma},$
which implies $(\phi,\psi)\in\tilde{\mathcal{G}}_{\omega,c,\alpha-3\gamma}$. ∎
@@
Let $1\leq N\leq 3$, $\gamma>0$ and $\gamma\neq 3$. Assume that $\tilde{a}_{\omega,c,\alpha-3\gamma}\neq\emptyset$, then $\widetilde{\mathcal{G}}_{\omega,c,\alpha-3\gamma}\subset\tilde{a}_{\omega,c,\alpha-3\gamma}$.
@@
Let $(\phi,\psi)\in\widetilde{\mathcal{G}}_{\omega,c,\alpha-3\gamma}$. One can take $(\phi_{1},\psi_{1})\in\tilde{a}_{\omega,c,\alpha-3\gamma}$, by Lemma 2.3, we have $(\phi_{1},\psi_{1})\in\widetilde{\mathcal{G}}_{\omega,c,\alpha-3\gamma}$, i.e., $\tilde{S}_{\omega,c,\alpha-3\gamma}(\phi_{1},\psi_{1})=\tilde{S}_{\omega,c,\alpha-3\gamma}(\phi,\psi)$. Therefore, for each $(u,v)\in\tilde{\mathcal{N}}_{\omega,c,\alpha-3\gamma}$, we obtain
$\tilde{S}_{\omega,c,\alpha-3\gamma}(\phi,\psi)=\tilde{S}_{\omega,c,\alpha-3\gamma}(\phi_{1},\psi_{1})\leq\tilde{S}_{\omega,c,\alpha-3\gamma}(u,v).$
Since $(\phi,\psi)\in\widetilde{\mathcal{G}}_{\omega,c,\alpha-3\gamma}\subset\tilde{\mathcal{N}}_{\omega,c,\alpha-3\gamma}$, we deduce that $(\phi,\psi)\in\tilde{a}_{\omega,c,\alpha-3\gamma}$. ∎
@@
Let $1\leq N\leq 3$, $\gamma>0$ and $\gamma\neq 3$. Then there exists $M>0$ such that $\tilde{\mu}_{\omega,c,\alpha-3\gamma}={\mu}_{\omega,c,\alpha-3\gamma}\geq\frac{1}{2M}$.
@@
By the definition of $\tilde{S}_{\omega,c,\alpha-3\gamma}$, $\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}$ and $\tilde{N}_{\omega,c,\alpha-3\gamma}$, we have the following relation
$\tilde{S}_{\omega,c,\alpha-3\gamma}=\frac{1}{2}\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}+\frac{1}{4}\tilde{N}_{\omega,c,\alpha-3\gamma}.$
One can rewrite $\tilde{\mu}_{\omega,c,\alpha-3\gamma}$ as
$\tilde{\mu}_{\omega,c,\alpha-3\gamma}=\inf\left\{\frac{1}{2}\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}:~{}~{}(\phi,\psi)\in\tilde{\mathcal{N}}_{\omega,c,\alpha-3\gamma}\right\}.$ (2.8)
By the Hölder inequality, Gagliardo-Nirenberg inequality and Sobolev embedding theorem, we have
$\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}(u,v)=2\widetilde{D}(u,v)$
@@
$\leq$ $M_{1}\|\nabla u\|_{L^{2}}^{N}\|u\|_{L^{2}}^{4-N}+M_{1}\|\nabla v\|_{L^{2}}^{N}\|v\|_{L^{2}}^{4-N}+M_{3}\|\nabla u\|_{L^{2}}^{\frac{3N}{4}}\|u\|_{L^{2}}^{\frac{3(4-N)}{4}}\|\nabla v\|_{L^{2}}^{\frac{N}{4}}\|v\|_{L^{2}}^{\frac{(4-N)}{4}}$
$\leq$ $M\tilde{\mathcal{Q}}^{2}_{\omega,c,\alpha-3\gamma}(u,v),$
where $M=M(N)$ is independent of $c$ and $\gamma$. Dividing both sides of the above by $\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}>0$, we obtain the positive lower bound
$\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}\geq\frac{1}{M}.$
This means that
$\mu_{\omega,c,\alpha-3\gamma}=\tilde{\mu}_{\omega,c,\alpha-3\gamma}\geq\frac{1}{2M}.$
This completes the proof of Lemma 2.5. ∎
@@
Let $1\leq N\leq 3$, $\gamma>0$ and $\gamma\neq 3$. If $(u,v)\in\tilde{X}_{\omega,c,\alpha-3\gamma}$ satisfies $\tilde{N}_{\omega,c,\alpha-3\gamma}(u,v)<0$, then $\frac{1}{2}\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}>\tilde{\mu}_{\omega,c,\alpha-3\gamma}$.
@@
If $\tilde{N}_{\omega,c,\alpha-3\gamma}(u,v)<0$, then $02\tilde{D}(u,v)>\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}(u,v)>$ 0. From this, we have
$\lambda_{0}:=\frac{\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}(u,v)}{2\tilde{D}(u,v)}\in(0,1)$
and $\tilde{N}_{\omega,c,\alpha-3\gamma}(\lambda_{0}u,\lambda_{0}v)=0$. By (2.8), we deduce
$\tilde{\mu}_{\omega,c,\alpha-3\gamma}\leq\frac{1}{2}\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}(\lambda_{0}u,\lambda_{0}v)=\frac{\lambda_{0}^{2}}{2}\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}(u,v)<\frac{1}{2}\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}(u,v).$
@@
Let $1\leq N\leq 3$, $\gamma>0$ and $\gamma\neq 3$. If the sequence $\{(u_{n},v_{n})\}$ weakly converges to $(u,v)\in\tilde{X}_{\omega,c,\alpha-3\gamma}$, then
$\tilde{D}(u_{n},v_{n})-\tilde{D}(u_{n}-u,v_{n}-v)\to\tilde{D}(u,v)~{}~{}\text{as}~{}~{}n\to\infty.$
@@
Note that $\tilde{Q}_{\omega,c,\alpha-3\gamma}$ and $\tilde{D}$ are invariant under
@@
$\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}(\tilde{\tau}_{y}(u,v))=\tilde{\mathcal{Q}}_{\omega,c,\alpha-3\gamma}(u,v),~{}~{}\tilde{D}(\tilde{\tau}_{y}(u,v))=\tilde{D}(u,v),$